40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 2 00:54:08 2017, Koji, Update, ASS, Xarm ASS restoration work 7x
    Reply  Fri Jun 2 01:22:50 2017, gautam, Update, ASS, ASS restoration work ASS_Y_recovery.pngASS_Repairs.zipOLs.png
       Reply  Wed Jun 7 14:11:49 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdfY_OL_calib_check.png
          Reply  Wed Jun 7 17:45:11 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdf
             Reply  Thu Jun 8 02:11:28 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation AnalogCheck.pdfAnalogCheck_zoom.pdfArmCtrl.pdf
                Reply  Mon Jun 12 17:42:39 2017, gautam, Update, ASS, ETMY Oplev Pentek board pulled out 
                   Reply  Wed Jun 14 18:15:06 2017, gautam, Update, ASS, ETMY Oplev restored 
                      Reply  Thu Jun 15 01:56:50 2017, gautam, Update, ASS, ETMY Oplev restored ITMY_OLspec.pdf
                         Reply  Tue Jun 27 16:11:28 2017, gautam, Update, Electronics, Coil whitening ETMYT_1182669422.mp4ITMY_whitening_studies.pdf
Message ID: 13052     Entry time: Thu Jun 8 02:11:28 2017     In reply to: 13051     Reply to this: 13060
Author: gautam 
Type: Update 
Category: ASS 
Subject: Y-arm coil driver electronics investigation 

Summary:

I investigated the analog electronics in the coil driver chain by using awggui to drive a given channel with Uniform noise between DC and 8kHz, with an overall gain of 1000 cts. This test was done for both ITMs and the BS. The Whitening/De-Whitening was off during the test. I measured the spectra in

  1. The digital domain (with DTT)
  2. At the output monitor of the AI board (with SR785)
  3. At the output of the coil driver board (with SR785)

Attachment #1 - There is good agreement between all 3 measurements. To convert the DTT spectrum to Vrms/rtHz, I multiplied the Y-axis by 10V / ( 2*sqrt(2) * 2^15 cts). Between DC and ~1kHz, the measured spectrum everywhere is flat, as expected given the test conditions. The AI filter response is also seen.

Attachment #2 - Zoomed in view of Attachment #1 (without the AI filter part).

*The DTT plots have been coarse-grained to keep the PDF file size managable. X (Y) axes are shared for all the plots in columns (rows).

 

Similar verification remains to be done for the ETMs, after which the test has to be repeated with the Whitening/DeWhitening engaged. But it's encouraging that things make sense so far (except perhaps the coil balancing can be better as suggested by the previous elog). 

 

I've left both arms locked. The Y-arm dither alignment is working well again, but for the X arm, the loops that actuate on the BS are still weird. Nothing obvious in the tests so far though.

GV 6pm 8 Jun 2017: I realized the X arm transmission was being monitored by the high-gain PD and not the QPD (which is how we usually run the ASS). The ASC mini screen suggested the transmitted beam was reasonably well centered on the X end QPD, and so I switched to this after which the X end dither alignment too converged. Possibly the beam was falling off the other PD, which is why the BS loops, which control the beam spot position on the ETM, were acting weirdly.

Quote:

will investigate the analog part of the drive electronics now.

 

Not related to this work:

I noticed the X-arm LSC servo was often hitting its limit - so I reduced the gain from 0.03 to 0.02. This reduced the control signal RMS, and re-acquiring lock at this lower gain wasn't a problem either. See attachment #3 (will be rotated later) for control signal spectra at this revised setting.

Attachment 1: AnalogCheck.pdf  469 kB  Uploaded Thu Jun 8 03:18:25 2017  | Hide | Hide all
AnalogCheck.pdf
Attachment 2: AnalogCheck_zoom.pdf  505 kB  Uploaded Thu Jun 8 03:19:00 2017  | Hide | Hide all
AnalogCheck_zoom.pdf
Attachment 3: ArmCtrl.pdf  38 kB  Uploaded Thu Jun 8 13:58:35 2017  | Hide | Hide all
ArmCtrl.pdf
ELOG V3.1.3-